OpenStudy (anonymous):

i need help setting up a table for d=rt, this is the problem two planes leave LA at the same time .one heads south to San Diego while the other heads north to San Francisco. the sandiego plane flies 50mph slower than the SanF plane.in 1/2 hr the planes are 275 mi apart.what are their speeds?

OpenStudy (anonymous):

put one rate at r, the other at r+50 and solve \[r + r+50 = 550\]

OpenStudy (anonymous):

i did then i divided by 50

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

i did but i have no idea what to do with the 1/2 hr and planes being 275 miles apart

OpenStudy (anonymous):

i thought since you had miles per hour better to work with one hour instead of half an hour. that is why i doubled it

OpenStudy (anonymous):

\]2r+50=550\] \[2r=500\] \[r=250\] and \[r+50=300\]

OpenStudy (anonymous):

that is good i am having trouble setting up the chart ..do you mind explaining that to me

OpenStudy (anonymous):

i do not know what a chart would look like

OpenStudy (anonymous):

ok, may i ask how did you get 550

OpenStudy (anonymous):

got it thank u